### v2

# The Baldwin Effect in Advancing Generalizability of Physics-Informed Neural Networks

Abstract
Physics-informed neural networks (PINNs) are at the forefront of scientific machine learning, making possible the creation of machine intelligence that is cognizant of physical laws and able to accurately simulate them. However, today’s PINNs are often trained for a single physics task and require computationally expensive re-training for each new task, even for tasks from similar physics domains. To address this limitation, this paper proposes a pioneering approach to advance the generalizability of PINNs through the framework of Baldwinian evolution. Drawing inspiration from the neurodevelopment of precocial species that have evolved to learn, predict and react quickly to their environment, we envision PINNs that are pre-wired with connection strengths inducing strong biases towards efficient learning of physics. A novel two-stage stochastic programming formulation coupling evolutionary selection pressure (based on proficiency over a distribution of physics tasks) with lifetime learning (to specialize on a sampled subset of those tasks) is proposed to instantiate the Baldwin effect. The evolved Baldwinian-PINNs demonstrate fast and physics-compliant prediction capabilities across a range of empirically challenging problem instances with more than an order of magnitude improvement in prediction accuracy at a fraction of the computation cost compared to state-of-the-art gradient-based meta-learning methods. For example, when solving the diffusion-reaction equation, a 70x improvement in accuracy was obtained while taking 700x less computational time. This paper thus marks a leap forward in the meta-learning of PINNs as generalizable physics solvers. Sample codes are available at https://github.com/chiuph/Baldwinian-PINN.
Index Terms:
Baldwin effect, neuroevolution, meta-learning, physics-informed neural networks
I Introduction
The emerging field of scientific machine learning seeks to create more accurate, data-efficient, and explainable machine intelligence for science and engineering. The direct incorporation of mathematically expressible laws of nature into learned models to ensure physically consistent predictions is an appealing proposition, as evidenced by the proliferation of physics-informed neural networks (PINNs) across multiple scientific domains since seminal work by Raissi et al. [1]. The key concept is to utilize physics-based mathematical relations or constraints as a regularization loss (aka physics-informed loss). This physics-informed loss is amenable to various forms of scientific knowledge and theories, including fundamental ordinary or partial differential equations (ODEs or PDEs). It flexibly incorporates scientific discoveries accumulated across centuries into the machine intelligence models of today across diverse scientific and engineering disciplines [2, 3, 4, 5, 6, 7, 8].
However, PINNs remain limited in their ability to generalize across physics scenarios. Contrary to its promise, a PINN does not guarantee compliance with physics when used for new scenarios unseen during training, e.g., variations in PDE parameters, initial conditions (ICs) or boundary conditions (BCs) that lie outside the confines of their training. Instead, these predictions remain physics-agnostic and may experience similar negative implications for reliability as typical data-driven models.
In principle, physics-compliant predictions for any new scenario can be achieved by performing physics-based retraining—an attractive feature of PINNs—even without labelled data. However, the additional training can be cost-prohibitive as physics-based learning is more difficult than data-driven learning due to the ruggedness of physics-informed loss landscapes, even with state-of-the-art gradient-based optimization algorithms [9, 10, 11, 12]. This has motivated the exploration of transfer learning techniques where connection strengths from similar (source) physics scenarios are used to facilitate accurate learning of solutions for new, harder problems [13, 14]. The related notion of meta-learning seeks to discover an optimized initialization of a model to enable rapid adaptation to a new test task with minimal training [15]. Nonetheless, most transfer- and meta-learned PINNs proposed to date still require the aforementioned physics-based retraining (with a substantial number of optimization iterations) to achieve more accurate solutions, and are therefore not ideal for applications that call for repeated, fast evaluations. A method to arrive at a generalizable PINN, one that can provide fast and accurate physics prediction/simulation across a varied set of unseen scenarios, remains elusive.
Figure 1: Schematic diagram of (a) Baldwinian evolution in nature and (b) evolving machine intelligence for learning physics with Baldwinian-PINNs. In nature, the Baldwin effect describes how learned traits are eventually reinforced in the genetic makeup of a population of organisms through natural selection. Equivalently, a population of Baldwinian-PINNs evolves over generations by being exposed to a broad distribution of physics tasks, gradually reinforcing traits promoting accurate physics learning into their genetic makeup. The evolved Baldwinian-PINNs are inherently equipped with strong learning biases to accurately solve any physics tasks over a broad task distribution.
In search of machine intelligence that generalizes for physically-consistent simulations of varied processes in the natural world, this paper studies the meta-learning of PINNs through the lens of neural Baldwinism—an expression of the Baldwin effect in the evolution of brains and intelligence [16]. Inspiration is drawn from the transmission of knowledge and predispositions across generations in precocial species, whereby their young are “born ready” with strong learning biases to perform a wide range of tasks. In order to pre-wire such learning ability into the initial connection strengths of a neural network, we examine an algorithmic realization of neural Baldwinism in the context of PINNs. The essence of Baldwinism lies in the phenomenon that characters learned by individual organisms of a group during their lifetime may eventually, under selection pressure, get reinforced by associated hereditary characters [17]. Analogously, Baldwinian neuroevolution of PINNs consists of an outer evolutionary optimization loop in which populations of PINN models are collectively exposed to a wide range of physics tasks sampled from a probability distribution over tasks of interest. Models with higher propensity to perform a random subset of those tasks well in an inner lifetime learning loop are evaluated as being fitter for survival, thus inducing a selection pressure towards connection strengths that encode stronger learning biases. The parallels between neural Baldwinism in nature and that of PINNs is depicted in Figure 1. The mathematical formulation of the problem resembles a two-stage stochastic program [18, 19], where the solution to lifetime learning enables the PINN to rapidly specialize to a specific physics task at test time.
Harnessing evolutionary procedures to optimize neural networks lends much greater versatility relative to other meta-learning approaches in terms of jointly crafting network architectures, initial network connection strengths, as well as learning algorithms, all through the use of potentially non-differentiable fitness functions [20]. Such biological “neuroevolution” [21, 22] precludes the need for explicit parameterization of tasks, facilitating generalization over task distributions comprising any broad mix of ODE/PDEs, ICs, and BCs. The physics-based lifetime learning of the neural network can be accelerated by reduction to a least-squares learning problem in its output layer, making Baldwinian neuroevolution computationally feasible. Such a least-squares formulation yields a closed-form result by means of the Moore-Penrose pseudoinverse and guarantees zero physics-informed loss given a sufficiently overparameterized network. The closed-form expression vastly reduces (or even eliminates) the need for iterative parameter updates to enable extremely fast lifetime learning of desired physics. Critically, the evolutionary search procedure is inherently highly parallelizable, thereby allowing for efficient meta-learning at scale by capitalizing on state-of-the-art advances in multi-CPU/GPU hardware infrastructure.
Neural Baldwinism thus makes it possible to achieve generalizable PINNs that are “genetically equipped” to perform well over a wide range of physics tasks. The evolved models (referred to as Baldwinian-PINNs) are demonstrated in this study to be broadly applicable to the simulation of families of linear and nonlinear ODE/PDEs encompassing diverse physical phenomena such as particle kinematics, heat and mass transfer, and reaction-diffusion. These Baldwinian-PINNs are capable of fast and accurate physics-aware predictions on previously unseen tasks, demonstrating up to several orders of magnitude computation speedup along with an order of magnitude improvement in prediction accuracy relative to recent meta-learned PINNs [15].
The key contributions of this paper are summarized below.
  • This is the first study to unveil Baldwinian evolution as a compelling route towards discovering neural nets with high capacity to learn diverse physics tasks, thereby advancing generalizability in PINNs.
  • An instantiation of the Baldwinian evolution framework is proposed through a novel two-stage stochastic programming formulation, wherein the first stage evolves the initial layers of a generalizable PINN model and the second stage trains its final layers to specialize to any new physics task (analogous to lifetime learning).
  • Comprehensive numerical experimentation and analysis shows that this methodology produces PINNs that effectively learn and predict across multiple PDE problems spanning different physics domains, demonstrating significant advancements in terms of speed and accuracy over models meta-learned by gradient descent.
The remainder of the paper is organized as follows. The basic problem setup for a PINN, the notion of generalization over physics tasks, and an overview of related work in the literature are presented in Section II. The proposed methodology for evolving Balwinian-PINNs is detailed in Section III. Extensive numerical assessment of the method is then carried out in Section IV over a range of linear and nonlinear ODE/PDEs. The paper is concluded in Section V with a discussion on directions for future research.
II Preliminaries
II-A Problem setup
II-A1 Single PINN problem
For simplicity of exposition, let us consider a problem with single spatial dimension $x$, time dimension $t$, and a single variable of interest $u$. In general, PINNs can learn a mapping between the input variables $(x,t)$ and the output variable $u$ while satisfying specified governing equations representing the physical phenomenon or dynamical process of interest:
PDE: $\mathcal{N}_{\_}\vartheta[u(x,t)]$ $=h(x,t),$ $x\in\Omega,t\in(0,T]$ (1a)
IC: $u(x,t=0)$ $=u_{\_}0(x),$ $x\in\Omega$ (1b)
BC: $\mathcal{B}[u(x,t)]$ $=g(x,t),$ $x\in\partial\Omega,t\in(0,T]$ (1c)
where the general differential operator $\mathcal{N}_{\_}\vartheta[u(x,t)]$ can include linear and/or nonlinear combinations of temporal and spatial derivatives and PDE parameters $\vartheta$, and $h(x,t)$ is an arbitrary source term in the domain $x\in\Omega,t\in(0,T]$. The IC (Eq. 1b) specifies the initial state, $u_{\_}0(x)$, at time $t=0$, and the BC (Eq. 1c) specifies that $\mathcal{B}[u(x,t)]$ equates to $g(x,t)$ at the domain boundary $\partial\Omega$.
Crucially, individual PINN models arrive at an accurate and physics-compliant prediction $u(x,t)$ for a single target scenario by minimizing the discrepancy between Eq. 1 and the model’s prediction during training.
II-A2 Generalizable neural physics solver
While most PINN models are trained to solve a specific physics task, there is increasing interest in generalizable neural physics solvers, i.e. models that can be flexibly applied to multiple physics problems once trained. In this context, we can consider a physical phenomena of interest that is represented by a set of training tasks belonging to some underlying task-distribution $p(\bm{\mathcal{T}})$, e.g., a family of PDEs spanning different PDE parameters, ICs and/or BCs.
Hence, the goal is to discover generalizable PINN models capable of fast, accurate, and physics-aware predictions on unseen scenarios, i.e., any new task from the distribution, $\mathcal{T}_{\_}i\backsim p(\bm{\mathcal{T}})$, by learning the underlying governing physics. In the context of meta-learning, the learning objective is to use training tasks from $p(\bm{\mathcal{T}})$ to find network initializations that are most amenable to a quick and accurate solution for the PINN loss, thereby accelerating the solution of multiple related physics problems at test time and providing a potential route to a generalizable neural physics solver.
II-B Related Work
PINN models are usually trained to solve a single, specific physics task. However, recent studies on meta-learning of PINNs to solve different physics scenarios as separate tasks have emerged, although no work has been reported from the neuroevolution perspective to our knowledge.
Most of the meta-learning PINN approaches reported in literature use weight interpolation as the basic framework. The simplest instantiation of this approach is to first train independent PINN models for each task and then interpolate across model weights for the new task [15, 23]. Several interpolation methods such as the Gaussian Process (GP) and Radial Basis Function (RBF) have been studied and shown to improve physics-informed learning on new tasks [15]. In other studies, interpolation is learned through the use of hypernetworks, whereby the hypernetworks and PINN are trained simultaneously from all the tasks [24, 25, 26]. Other variants include encoding tasks as latent variables and passing them into the input layer of the PINN model [27].
However, there are two major drawbacks to such weight interpolation frameworks. Firstly, these methods rely heavily on task parameterization to perform interpolation and operate under the assumption of smoothness. This means that all the train tasks (and the new task) must adhere to this parameterization requirement. In addition, these methods do not capitalize on the principal characteristic of PINN, which is the potential for physics-informed learning (retraining or fine-tuning) for a new task, in their meta-learning formulation. The data-free nature of physics-informed fine-tuning is not exploited during the meta-learning phase, given that there is no guiding principle on how the fine-tuning towards any new task should be performed given the interpolated weights.
Figure 2: (a) Meta-learning PINN with Baldwinian neuroevolution (right) versus MAML (left). In MAML, the initial weights ${\bf\it W}$ are learned using gradient-based method, such that they can be quickly fine-tuned (physics-informed learning) on new tasks. Although the task-specific fine-tuning is limited to one or a few gradient descent updates during training, such amount of fine-tuning is usually insufficient for a PINN at test time. In Baldwinian neuroevolution, the weight distribution in the pre-final nonlinear hidden layers ${\bf\it\tilde{w}}$ and learning hyperparameters $\lambda$ are jointly evolved. The task-specific physics-informed learning is performed on the output layer weights $w$ (segregated from ${\bf\it\tilde{w}}$) with a 1-step pseudoinverse operation at both training and test time (for linear ODE/PDEs). (b) Schematic of Baldwinian-PINNs architecture used in present study and procedure to obtain nonlinear hidden layers’ weights ${\bf\it\tilde{w}}$’s from the evolved network hyperparameter $\theta$ for learning task-specific outputs.
The model-agnostic meta-learning (MAML) [28] framework can theoretically overcome the limitations of weight interpolation methods. MAML and its variant, Reptile [29], aim to learn an optimal weight initialization that can be quickly fine-tuned on new tasks without the need for interpolation [30]. An illustration of MAML for PINN is shown in Figure 2(a). To reduce the computation cost of meta-learning, task-specific physics-informed learning is limited to one or a few gradient descent updates from the initialization. This essentially assumes that the model performance after one step (or a few steps of) gradient descent update from the initialization is already indicative of the learning performance. While this assumption may be appropriate in data-driven models, PINN models are much more challenging to optimize. Hence, they require more thorough training (e.g., a large number of updates, even with the use of 2nd-order gradient descent methods) to achieve a good convergence. Moreover, gradient-based MAML methods are prone to getting stuck in a local minimum and struggle to find a good initialization for a diverse set of tasks, as the gradients can be noisy or even deceptive, i.e., requiring going against it to reach the optimum. Hence, recent studies have shown poor MAML PINN results, potentially due to non-convergence during meta-learning or insufficient physics-informed fine-tuning performed at test time [15, 24].
III Proposed Methodology
We address the above limitations with a novel meta-learning PINN framework based on Baldwinian neuroevolution. Baldwinian neuroevolution aims to optimize the weight distribution in the pre-final nonlinear hidden layers and physics-informed learning hyperparameters that are essential for achieving optimal performance in task-specific physics-informed learning. We limit our task-specific learning to the output layer, allowing us to solve linear problems with a 1-step pseudoinverse operation (or a few steps for nonlinear problems) during both the meta-learning process and test time, in a consistent manner. Importantly, once given the nonlinear hidden layers’ weights and learning hyperparameters, the pseudoinverse operation is exact and fast (does not require further gradient descent updates). An illustration of Baldwinian neuroevolution for PINN is shown in Figure 2(a).
III-A A two-stage stochastic programming formulation
Without loss of architectural generality, Baldwinian-PINNs can be assigned the form of a multilayer perceptron (MLP) with proven representation capacity [31, 32, 33], thereby ensuring the ability to learn a diversity of tasks. Its output $u(x,t)$ can be written as:
$u(x,t)=\sum_{\_}jw_{\_}jf_{\_}j(x,t;{\bf\it\tilde{w}})$ (2)
where ${\bf\it w}=[\dots w_{\_}j\dots]^{T}$ are the output layer weights, and $f_{\_}j(x,t;{\bf\it\tilde{w}})$’s represent nonlinear projections of the input variables with the hidden layers’ weights ${\bf\it\tilde{w}}$. The connection strengths ${\bf\it\tilde{w}}$ are typically learned during PINN training. However, in the proposed implementation of Baldwinian-PINN as detailed below, they are drawn from a probability distribution defined by a dimensionally reduced set of network hyperparameters $\theta$ (see example in Section III-B), that are assigned at birth as per Baldwinian evolution.
The model’s reaction to a given environment—i.e., finding the best set of $w_{\_}j$’s such that the Baldwinian-PINN’s output satisfies Eq. 1 for a specific physics task $\mathcal{T}_{\_}i$—is the focus of a typical single PINN problem described in Section II-A. This is reduced to a physics-based least-squares problem in this work:
$\operatorname*{arg\,min}_{\_}{{\bf\it w}}\ \mathbf{(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}-\mathbf{b}^{\mathcal{T}_{\_}i})^{T}}(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}-\mathbf{b}^{\mathcal{T}_{\_}i})+\lambda{\bf\it w}^{\mathbf{T}}{\bf\it w}$ (3)
where $\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}$ is obtained by substituting the model’s output into the left hand side of Eq. 1 for a given set of collocation points, and $\mathbf{b}^{\mathcal{T}_{\_}i}$ represents the corresponding right hand side of Eq. 1. This is chosen as Eq. 3 yields a closed-form result when using the Moore-Penrose generalized inverse, permitting extremely fast learning (order of milli-seconds in our experiments) for physics-compliant prediction. Similar least-squares formulations studied in the literature have been shown to be competitive with, or faster than, widely-used numerical solvers such as the finite element method [34, 35]. The learning hyperparameter $\lambda\geq 0$ reduces the $L^{2}$-norm of the least-squares solution, thereby improving the solution numerically. Detailed derivations, including information on the construction of matrix $\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}$ and vector $\mathbf{b}^{\mathcal{T}_{\_}i}$, are provided in Section III-C.
Going beyond specialization to a single physics task by means of Eq. 3, we mathematically formulate the search for Baldwinian-PINN models that can generalize to a whole family of PDEs as the following two-stage stochastic optimization problem [19]:
$\min_{\_}{\theta,\lambda}\operatorname{\mathbb{E}}_{\_}{\mathcal{T}_{\_}i\backsim p(\bm{\mathcal{T}})}\operatorname{\mathbb{E}}_{\_}{{\bf\it\tilde{w}}\backsim p_{\_}{\theta}({\bf\it\tilde{w}})}$ $[\tau_{\_}{LSE}\ l_{\_}{LSE}({\bf\it w^{*}})+\tau_{\_}{MSE}\ l_{\_}{MSE}({\bf\it w^{*}})]\quad$
$\text{subject to }\tau_{\_}{LSE}\geq 0,\ \tau_{\_}{MSE}\geq 0$ (4)
where the solution ${\bf\it w^{*}}$ of the second stage problem:
${\bf\it w^{*}}=\operatorname*{arg\,min}_{\_}{{\bf\it w}}\ \mathbf{(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}-\mathbf{b}^{\mathcal{T}_{\_}i})^{T}}(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w}-\mathbf{b}^{\mathcal{T}_{\_}i})+\lambda{\bf\it w}^{\mathbf{T}}{\bf\it w}$ (5)
provides optimal output layer weights that allow the model to specialize to any realization of task $\mathcal{T}_{\_}i\backsim p(\bm{\mathcal{T}})$ for the given network’s ${\bf\it\tilde{w}}\backsim p_{\_}{\theta}({\bf\it\tilde{w}})$.
The first stage optimization objective is defined by the weighted sum of the physics learning proficiency, i.e., sum of squared residuals or least-squares error (LSE),
$l_{\_}{LSE}({\bf\it w^{*}})=\mathbf{(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w^{*}}-\mathbf{b}^{\mathcal{T}_{\_}i})^{T}}(\mathbf{A}_{\_}{{\bf\it\tilde{w}}}^{\mathcal{T}_{\_}i}{\bf\it w^{*}}-\mathbf{b}^{\mathcal{T}_{\_}i})$ (6)
and the actual predictive performance, i.e., mean squared error (MSE),
$l_{\_}{MSE}({\bf\it w^{*}})=\frac{1}{n}\sum_{\_}{s=1}^{n}\ (u_{\_}s^{label}-\sum_{\_}jw_{\_}j^{*}f_{\_}j(x_{\_}s,t_{\_}s;{\bf\it\tilde{w}}))^{2}$ (7)
given labelled data $u_{\_}s^{label},s=1,...,n$, over the task distribution $p(\bm{\mathcal{T}})$ and the network connections’ distribution $p_{\_}{\theta}({\bf\it\tilde{w}})$.
III-B Baldwinian-PINN instantiation: shallow network architecture and connections at birth
In the present study, Baldwinian-PINNs are designed to have weights and biases in the nonlinear hidden layers fixed to random values at birth—a randomized neural networks setup [36, 37]—drawn from a probability distribution defined by a dimensionally reduced set of network hyperparameters $\theta$. Both normal and uniform distributions for the weights and biases are possible for the randomized PINNs [34, 35]. Similarly, smooth nonlinear activations such as sin, softplus, and tanh are common in PINN literature, and can be advantageous for different problems. For greater flexibility, we apply both distributions for setting weights and biases and all three activations to different hidden layer blocks in the Baldwinian-PINN models.
Hence, the base Baldwinian-PINNs use a single hidden layer architecture which is segmented into $3\times 2=6$ unique blocks, with the weights and biases from 3 blocks sampling from normal distributions and the weights and biases from another 3 blocks sampling from uniform distributions. Each block has a fixed number of neurons ($n_{\_}{neuron}$ = 150 or 200). Assuming 2 input variables $(x,t)$, we can write the output $f_{\_}j$ for all the neurons $j=1,...,n_{\_}{neuron}$ in a nonlinear hidden-layer block as:
$y_{\_}j=\tilde{w}_{\_}j^{3b-2}x+\tilde{w}_{\_}j^{3b-1}t+\tilde{w}_{\_}j^{3b}$ (8a)
$f_{\_}j(x,t;{\bf\it\tilde{w}}_{\_}j)=\varphi^{b}(y_{\_}j)$ (8b)
for each of the blocks $b=1,...,6$, where the activation $\varphi^{b}$ can be sin ($b=1,4$), softplus ($b=2,5$), or tanh ($b=3,6$). The $\tilde{w}_{\_}j^{i}$’s are weights and biases with their own distributional mean $m^{i}$ and spread $s^{i}$, $i=1,...,18$. Their values can be obtained by the following sampling procedure:
$\tilde{w}_{\_}j^{i}\sim\mathcal{N}(0,1)\quad\text{or}\quad\tilde{w}_{\_}j^{i}\sim\mathcal{U}(-1,1)$ (9a)
$\tilde{w}_{\_}j^{i}\leftarrow\tilde{w}_{\_}j^{i}\times s^{i}+m^{i}$ (9b)
The base Baldwinian-PINN model is depicted in Figure 2(b). Given the Baldwinian-PINNs’ configuration, a set of network hyperparameters $\theta=(m^{1},s^{1},...,m^{18},s^{18})$ control the distributional mean and spread of the weights and biases in different blocks. In the spirit of the Baldwin effect, the task-specific output layer ${\bf\it w}$ is the outcome of lifetime learning and not directly inherited by the next generation of offspring [38]; only the hyperparameters $(\theta,\lambda)$ are subjected to evolutionary variation and inheritance. Since the neuroevolution only searches the distribution parameters for groups of weights instead of evolving each individual weight in the nonlinear hidden layer, the reduced dimensionality can be even more effectively searched by today’s evolutionary algorithms.
While the above description is for a single hidden layer, we note that the Baldwinian-PINN framework is not restricted to a single hidden layer. We include additional examples in Suppl. LABEL:subsec:additional_evo_depth to show that the methodology also works for deeper neural architectures.
III-C Baldwinian-PINN lifetime learning
The lifetime learning procedure of the Baldwinian-PINNs occurs only in the linear output layer of the network, i.e., finding the best set of $w_{\_}j$’s such that the output $u(x,t)=\sum_{\_}jw_{\_}jf_{\_}j(x,t;{\bf\it\tilde{w}})$ satisfies the governing equations in Eq. 1 for a specific task. Given a set of collocation points $(x_{\_}i^{pde},t_{\_}i^{pde}),i=1,...,n_{\_}{pde},(x_{\_}i^{ic},0),i=1,...,n_{\_}{ic},(x_{\_}i^{bc},t_{\_}i^{bc}),i=1,...,n_{\_}{bc}$ sampled from the respective domain, the following system of equations can be formed:
$\mathbf{A}{\bf\it w}$ $=\mathbf{b}$ (36)
Note that the derivatives required to construct $\mathbf{A}$ can be easily computed by automatic differentiation [39]. The best-fit solution to the above system of linear equations with unknown ${\bf\it w}=[\dots w_{\_}j\dots]^{T}$ can be obtained by means of the Moore-Penrose pseudoinverse:
${\bf\it w}$ $=(\lambda I+\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}\mathbf{b}$ (37a)
(if the system of equations in Eq. 36 is over-determined)
${\bf\it w}$ $=\mathbf{A}^{T}(\lambda I+\mathbf{A}\mathbf{A}^{T})^{-1}\mathbf{b}$ (37b)
(if the system of equations in Eq. 36 is under-determined)
where the learning hyperparameter $\lambda\geq 0$ reduces the $L^{2}$-norm of the least-squares solution, thereby improving the solution numerically. Such a least-squares formulation yields a closed-form result in a single computation step for any linear PDE (i.e., the governing equations in Eq. 1 are all linear with respect to $u$) which encapsulates a wide class of problems in the natural world.
For nonlinear PDEs, iterative methods can be used for arriving at optimized $w_{\_}j$’s in Baldwinian-PINNs. In this work, we use a lagging of coefficient approach which is common in numerical methods [40] and detailed in Suppl. LABEL:subsec:lagging. Briefly, we approximately linearize the nonlinear term(s) in Eq. 1 by substituting the output $u(x,t)=\sum_{\_}jw_{\_}jf_{\_}j(x,t;{\bf\it\tilde{w}})$ obtained from previous step, and iteratively solve Eq. 37 to update $w_{\_}j$’s for a fixed number of steps, $N$, or until a convergence criterion is reached.
Hence, the Baldwinian-PINNs’ lifetime learning is broadly applicable to both linear and nonlinear PDEs, with the psuedoinverse formulation permitting extremely fast computation for physics-compliant prediction. In addition, there is no requirement on prior labelled data for any new task as the learning can be entirely physics-based (as per conventional PINNs).
III-D Baldwinian neuroevolution algorithm and hyperparameters
Algorithm 1 and Algorithm 2 describe a Baldwinian neuroevolution procedure to solve the two-stage stochastic programming problem defined in Section III-A. Recall a Baldwinian-PINN individual can be represented by $(\theta,\lambda)$, i.e., network architecture and learning algorithm hyperparameters, during neuroevolution.
Algorithm 1 Baldwinian neuroevolution of PINNs
1:$p(\bm{\mathcal{T}})$: distribution over physics tasks
2:$\bm{\mathcal{P}}$: initial population-representation of individuals (i.e., network architecture and learning algorithm hyperparameters in Baldwinian-PINN models)
3:$\bm{\mathcal{S}}$: procedure to obtain a batch of individuals given a population-representation
4:$\bm{\mathcal{U}}$: procedure to update a population-representation given a batch of fitness-scored individuals
$\bm{\mathcal{F}}$: procedure to return fitness based on lifetime learning performance of individual given a set of tasks
6:while not done do
7:     Sample batch of individuals from population $(\theta^{g},\lambda^{g})\backsim\bm{\mathcal{S}}(\bm{\mathcal{P}}),g=1,...,n_{\_}{pop}$
8:     Sample batch of tasks $\mathcal{T}_{\_}i\backsim p(\bm{\mathcal{T}}),i=1,...,n_{\_}{task}$
9:     for all $(\theta^{g},\lambda^{g})$ do
[…]
11:     end for
12:     Update population based on fitness of individuals:
13:      $\bm{\mathcal{P}}\leftarrow\bm{\mathcal{U}}(\bm{\mathcal{P}},\{(\theta^{g},\lambda^{g},f^{g}),g=1,....,n_{\_}{pop}\})$
14:end while
Algorithm 2 Baldwinian-PINN lifetime learning and fitness calculation $\bm{\mathcal{F}}$
INPUT: network architecture and learning algorithm hyperparameters $(\theta,\lambda)$, list of tasks $\mathcal{T}_{\_}i,i=1,...,n_{\_}{task}$
OUTPUT: fitness $f$
1:$\bm{\mathcal{G}}$: procedure to sample weights and biases in nonlinear hidden-layer(s)
2:$\bm{\mathcal{C}}$: procedure to construct least squares problem on a set of collocation points based on underlying physics (PDEs, BCs, ICs) of the task
3:$N$: number of iterations to take during per-task “lifetime learning”: $1$ for linear PDEs, $>1$ for nonlinear PDEs
4:Sample hidden layers’ weights and biases ${\bf\it\tilde{w}}=\bm{\mathcal{G}}(\theta)$
5:for all $\mathcal{T}_{\_}i$ do
6:     ${\bf\it w}=[\dots w_{\_}j\dots]^{T}=\mathbf{0}$triangleright initial solution for nonlinear PDEs
7:     for $k=1,...,N$ do
8:         Construct least squares matrix and vector $(\mathbf{A},\mathbf{b})=\bm{\mathcal{C}}(\mathcal{T}_{\_}i,{\bf\it\tilde{w}},{\bf\it w})$
9:         Compute least-squares solution via pseudoinverse:
10:            ${\bf\it w}=(\lambda I+\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}\mathbf{b}$ ,
11:            if $\mathbf{A}$ has more rows than columns
12:            ${\bf\it w}=\mathbf{A}^{T}(\lambda I+\mathbf{A}\mathbf{A}^{T})^{-1}\mathbf{b}$ ,
13:            if $\mathbf{A}$ has more columns than rows
14:     end for
15:     ${\bf\it w}^{\mathcal{T}_{\_}i}\leftarrow{\bf\it w}$, $\mathbf{A}^{\mathcal{T}_{\_}i}\leftarrow\mathbf{A}$, $\mathbf{b}^{\mathcal{T}_{\_}i}\leftarrow\mathbf{b}$
16:end for
[…]
18:Compute mean squared error: $l_{\_}{MSE}=\sum_{\_}{\mathcal{T}_{\_}i}\frac{1}{n}\sum_{\_}s^{n}(u_{\_}s^{label}-\sum_{\_}jw_{\_}j^{\mathcal{T}_{\_}i}f_{\_}j(x_{\_}s,t_{\_}s;{\bf\it\tilde{w}}_{\_}j))^{2}$ over the labelled data $u_{\_}s^{label},s=1,...,n$
19:Compute overall fitness: $f=-(\tau_{\_}{LSE}\ l_{\_}{LSE}+\tau_{\_}{MSE}\ l_{\_}{MSE})$
The Baldwinian neuroevolution procedure described in Algorithm 1 is generic for neuroevolution methods, such as covariance matrix adaptation evolution strategy (CMA-ES) [41]. As an instantiation of information-geometric optimization algorithms [42], CMA-ES represents the population of $(\theta,\lambda)$ in $\bm{\mathcal{P}}$ using a multivariate normal search distribution, initialized with zeros mean and standard deviation (std.) as tuning hyperparameter. $\bm{\mathcal{S}}$ becomes a random sampling procedure with population size or the number of individuals $n_{\_}{pop}$ to be sampled from the search distribution as hyperparameter, and $\bm{\mathcal{U}}$ is the CMA-ES update-step [41] which iteratively adapts its search distribution based on the rank-based fitness landscape until the convergence criteria, e.g., a pre-determined number of iterations or fitness value (trade-off between computation resource and convergence) is met.
The network weights and biases can take any value from (($\infty$( - ∞, $\infty$)∞ ), hence there is no restriction to the continuous search space of $\theta=(m^{1},s^{1},...)$ representing their distributional means and spreads. The learning hyperparameter $\lambda\geq 0$ can be evolved in continuous search space and its absolute value is then used for computing the least-squares solution. In our implementation with CMA-ES, $(\theta,\lambda)$ share the same initial standard deviation, and we scale the learning hyperparameter by a factor, i.e., $\lambda\leftarrow 1\mathrm{e}{-4}\times\text{abs}(\lambda)$, to improve the performance of the pseudoinverse.
Our preliminary experiments show that the performance of Baldwinian neuroevolution is robust across a range of CMA-ES hyperparameters such as population size and initial standard deviation (std.) of search distribution, and the number of tasks sampled for fitness evaluation per iteration. Hence, a robust setting that shows good convergence in fitness across different types of problems is chosen based on initial experiments. In the majority of this study, we employ the CME-ES for evolving $(\theta,\lambda)$, although other results in Suppl. LABEL:subsec:additional_evo_depth show the extensibility of this framework to other neuroevolution algorithms.
In Algorithm 2, a fixed set of collocation points are used for the construction of least squares matrix and vector $(\mathbf{A},\mathbf{b})$ across the training tasks for ease of implementation. We choose the collocation points to coincide with the location of labelled data for $l_{\_}{MSE}$ computation, although this is not a prerequisite. Similarly, the same number of nonlinear iterations $N$ is used across nonlinear training tasks sampled from the task distribution for ease of implementation.
In our preliminary experiments, we found it is critical to have the $l_{\_}{MSE}$ component in the optimization objective even though the learning of future test tasks remains solely physics-based. We set $\tau_{\_}{LSE}=\tau_{\_}{MSE}=1$ as default for the computation of overall fitness unless there is a huge difference in magnitude between $l_{\_}{LSE}$ and $l_{\_}{MSE}$ during Baldwinian neuroevolution. In addition, we can multiply the BC/IC rows in both $\mathbf{A}$ and $\mathbf{b}$ to re-balance the importance between PDE and BC/IC errors in the least-squares solution.
Although this study focuses on the paradigm of Baldwinian neuroevolution as a pathway towards generalizable neural physics solvers, other combinations of neuroevolution or lifetime learning algorithms may be used. One key advantage of neuroevolution is that the fitness evaluations (population size $n_{\_}{pop}$$\times$ number of random tasks $n_{\_}{task}$) required each iteration can be easily parallelized across multiple GPUs to fully harvest any hardware advantage. In particular, we utilized the JAX framework to harness previously reported performance improvements for automatic differentiation and linear algebra operations [43, 44]. The experimental study is performed on a workstation with an Intel Xeon W-2275 Processor and 2 NVIDIA GeForce RTX 3090.
IV Experimental Studies
We examine the efficacy of Baldwinian neuroevolution for learning physics (Algorithm 1 and Algorithm 2 in Section III-D) as formalized in a two-stage stochastic programming problem in  III-A. Several ODE/PDE problems which are representative of real-world phenomena are used to demonstrate Baldwinian neuroevolution for physics in the following sections. Table I summarizes various neuroevolution and Baldwinian-PINN lifetime learning configurations and performance on their respective test tasks.
IV-A Learning to solve and generalize linear ODE/PDEs
Figure 3: (a) Solution of 20 individual Baldwinian-PINN models sampled from the CMA-ES search distribution (initial std. = 1), for unseen convection-diffusion tasks $\alpha=\{4,27,108\}$. The task is more challenging with increasing $\alpha$. Baldwinian neuroevolution is effective for evolving good Baldwinian-PINN models which can generalize across different difficulties. (b) Baldwinian neuroevolution demonstrates effective LSE and MSE convergences on convection-diffusion problem, for different CMA-ES initial std. values (best std. = 1). The bold lines indicate their median convergence path from 5 individual runs, and the shaded areas indicate their interquartile ranges.
IV-A1 Convection-diffusion
The steady-state convection-diffusion equation is a ubiquitous physics model that describes the final distribution of a scalar quantity (e.g. mass, energy, or temperature) in the presence of convective transport and diffusion [45]. Solutions to this physics are key to characterization and design of many systems, including microfluidic chip cooling in electronics [46]. The 1D equation is defined as:
[…]
subject to BCs $u(x=0)=0$; $u(x=1)=1$. These real-world problems have characteristic physics that vary with non-dimensional constants such as the Peclet number $Pe$ (ratio of convection to diffusion) [47]. Hence, it is helpful to learn a PINN model that can return $u(x)$ for a diverse range of $Pe$-related problems (determined by $\alpha$ here). The training tasks consist of $\alpha=\{5,10,...,100\}$, encompassing both smoother output patterns at lower $\alpha$ and very high gradient patterns at higher $\alpha$, with the latter being challenging for PINNs to learn by both stochastic gradient descent (SGD) [13, 9] and classical numerical methods [48].
[…]
IV-A2 Family of linear PDEs (convection, diffusion, and dispersion)
Figure 4: (a) Schematic to illustrate new tasks arising from family of linear PDEs problem: S1 change to new PDE and IC profile for $t\in[0,2]$ (same time domain as train tasks), and S2 projection to longer time domain $t\in[0,4]$. (b) Solution for unseen linear PDE tasks obtained by best evolved Baldwinian-PINN sampled from the center of CMA-ES search distribution after 500 iterations with initial std. = 5 (they are visually indistinguishable from the ground truth). (c) The mean MSE over $n=87$ test tasks for 2 test scenarios described in (a) are below $5\mathrm{e}{-5}$. (d) Baldwinian neuroevolution demonstrates effective LSE and MSE convergences for different CMA-ES initial std. values, with a superior performance given by std. = 5 and 10. The bold lines indicate their median convergence path from 5 individual runs, and the shaded areas indicate their interquartile ranges.
Next, we extend Baldwinian-PINNs to a family of linear PDEs. This PDE family is a further generalization of the convection-diffusion equation:
$\text{(Problem 2)}\quad\frac{du}{dt}+\alpha\frac{du}{dx}-\gamma\frac{d^{2}u}{dx^{2}}+\delta\frac{d^{3}u}{dx^{3}}=q(x,t),\\ x\in[0,1],\quad t\in(0,2]$ (39a)
$q(x,t)=\sum_{\_}{j=1}^{J}A_{\_}j\sin\left(\omega_{\_}jt+\frac{2\pi l_{\_}jx}{L}+\varphi_{\_}j\right)$ (39b)
with IC $u(x,0)=q(x,0)$ and periodic BC $u(0,t)=u(1,t)$. Eq. 39 models the time evolution of a scalar quantity in the presence of physics phenomena such as convection ($\frac{du}{dx}$ component), diffusion ($\frac{d^{2}u}{dx^{2}}$ component), and dispersion ($\frac{d^{3}u}{dx^{3}}$ component), and a rich diversity of dynamical processes can be generated from different PDE and IC combinations [49, 50]. $q(x,t)$ is a source term, with different $q(x,t=0)$ being the corresponding IC profiles. We consider the following PDE scenarios: $\alpha=1$, $\gamma=\{0,5\mathrm{e}{-4},1\mathrm{e}{-3}\}$, $\delta=\{0,5\mathrm{e}{-4},1\mathrm{e}{-3}\}$. The ratio between $\alpha$, $\gamma$, and $\delta$ determine non-dimensional constants (e.g. $Pe$), and consequently, the systems’ characteristic physics. $q(x,t)$ comprises scenarios with $J=5$, $L=6$ and coefficients sampled uniformly from $A_{\_}j\in[-0.8,0.8]$, $\omega_{\_}j\in[-2,2]$, $l_{\_}j\in[0,1,2,3,4]$, $\varphi_{\_}j\in[-\pi,\pi]$. The training set comprises 108 tasks with different PDE and IC combinations.
[…]
Interestingly, the Baldwinian-PINNs maintain good accuracy on tasks from S2, whereby the evolved Baldwinian-PINN model first predicts $u(x,t)$ for $t\in[0,2]$,before using the solution $u(x,t=2)$ as new IC for $t\in[2,4]$. This is achievable because Baldwinian neuroevolution does not require parameterization for the tasks, and Baldwinian-PINNs can generalize to new ICs, BCs, and PDE source terms. This is tricky for existing meta-PINN methods as they require interpolation across a potentially infinitely large distribution of tasks (e.g. possible BCs or ICs), in contrast to the lifetime learning encapsulated in the Baldwinian paradigm. Additional results illustrating the ability of Baldwinian-PINNs to perform well despite variations in the lifetime learning task objectives (e.g., solving for different time durations) are in Suppl. LABEL:subsec:additional_linearpde_time, further emphasizing the merits of Baldwinian neuroevolution for physics.
Figure 5: (a) Solution of 20 individual Baldwinian-PINN models sampled from the CMA-ES search distribution (initial std. = 0.5), for unseen kinematics tasks. (b) Solution of Baldwinian-PINN model sampled from the center of CMA-ES search distribution after 100 iterations (initial std. = 1), for unseen 6D diffusion-reaction task. The solutions shown in (a) and (b) are visually indistinguishable from the ground truth. Baldwinian neuroevolution demonstrates effective LSE and MSE convergence on both (c) kinematics and (d) 6D parametric diffusion-reaction problem, for different CMA-ES initial std. values. The bold lines indicate their median convergence path from 5 individual runs, and the shaded areas indicate their interquartile ranges.
IV-A3 Additional linear ODE/PDE problems
The Baldwinian neuroevolution of physics is further demonstrated on 3 linear problems (Problems 3-5), namely 1D Poisson’s equation, 2D Poisson’s equation, and Helmholtz equation (see Suppl. LABEL:subsec:otherlinearpde2). Suppl. Table LABEL:tab:linearbenchmark enumerates the accuracy advantages of Baldwinian-PINN relative to results reported by recent meta-learning PINN works [30, 24]. Suppl. LABEL:subsec:otherlinearpde2 also presents additional visualization results from Baldwinian-PINN and other instantiations of the proposed Baldwinian neuroevolution framework, demonstrating the Baldwinian-PINNs’ versatility and generalizability.
IV-B Learning to solve and generalize nonlinear ODE/PDEs
IV-B1 Kinematics
Extending beyond linear ODE/PDEs, Baldwinian-PINNs are applied to nonlinear kinematics equations. Assuming a ball is thrown at specific launch angle $a_{\_}0$ and initial velocity $vel_{\_}0$, the following 2D kinematics equations describe the ball’s motion under the influence of gravity $g$ and air resistance $R$:
$\text{(Problem 6)}\quad\frac{d^{2}x}{dt^{2}}+R\frac{dx}{dt}=0,\\ t\in(0,T]$ (40a)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \frac{d^{2}y}{dt^{2}}+R\frac{dy}{dt}=-g,\\ t\in(0,T]$ (40b)
subject to ICs $x(t=0)=0,\frac{dx}{dt}(t=0)=vel_{\_}0\times\mathrm{cos}(\frac{a_{\_}0\pi}{180})$; $y(t=0)=0,\frac{dy}{dt}(t=0)=vel_{\_}0\times\mathrm{sin}(\frac{a_{\_}0\pi}{180})$. The air resistance $R=\frac{1}{2}\frac{\rho C_{\_}dA}{m}V$ is related to air density $\rho$, object properties (drag coefficient $C_{\_}d$, cross-sectional area $A$, and mass $m$), and object velocity $V=\sqrt{(dx\mathrm{/}dt)^{2}+(dy\mathrm{/}dt)^{2}}$, hence the equations are nonlinear with respect to $x$ and $y$. The 150 training tasks comprise different launch angles $a_{\_}0\in[15,85]$, initial velocity $vel_{\_}0\in[10,110]$, and object properties $C_{\_}d\in[0.2,0.7]$, $A\in[0.00145,0.045]$, $m\in[0.046,0.6]$ as may be representative of different projectiles (e.g., baseball, basketball). $g$ and $\rho$ are assumed to be $9.8$ and $1.3$ respectively.
[…]
IV-B2 Set of nonlinear PDEs
The Baldwinian neuroevolution of physics is further demonstrated on 5 nonlinear PDE problems (Problems 7-11) such as the Burger’s equation, nonlinear Allen-Cahn equation and nonlinear reaction-diffusion equation, as per recent meta-learning PINN study [15] and described in Suppl. LABEL:subsec:nonlinearpde2. That study compares several meta-learning PINNs based on weight interpolation methods and MAML.
A key difference between this study and meta-learning of PINNs in [15] is the availability of data for training tasks. On the test tasks, the evolved Baldwinian-PINNs can achieve $\backsim 1$ order of magnitude lower relative norm error relative to [15]. Crucially, the computation time for Baldwinian-PINN for new predictions is at most two seconds whereas other meta-learning approaches may take more than 500 seconds (2 orders of magnitude acceleration). For example, in numerical experiments involving the diffusion-reaction equation, a 70x improvement in accuracy was obtained relative to other state-of-the-art approaches, while taking 700x less computational time. Complete quantitative results are summarized in Suppl. Table LABEL:tab:benchmark. Representative Baldwinian-PINN results on 6D parametric diffusion-reaction problem are presented in Figure 5c-d.
The versatility of Baldwinian neuroevolution is also key here as we can accelerate the Baldwinian-PINN’s lifetime learning by using a much coarser discretization and smaller number of nonlinear iterations during training while still learning a good solution on test tasks with finer (e.g., 16$\times$) discretization and more (e.g., 2$\times$) nonlinear iterations.
Figure 6: The generalization performance of the DNN, PINN, and Baldwinian-PINN models are compared, across (a) 110 convection-diffusion and (b) 100 kinematics test tasks. Model-1a/2a: deep/shallow architecture with tanh activations (network weights initialized by Xavier method); Model-1b/2b: deep/shallow architecture with sin activations (weights initialized by He method). The MSE results from each of the DNN and PINN models are pooled from 3 initial learning rate configurations $\times$ 5 individual runs. The MSE results from Baldwinian-PINN model are pooled from 5 initial std. values $\times$ 5 individual runs.
IV-C Analysis on effectiveness of Baldwinian neuroevolution
We further investigate the advantages of Baldwinian neuroevolution via an ablation study using the convection-diffusion and kinematics examples.
Briefly, we explore deep and shallow MLP architectures for baseline SGD-trained DNN and PINN models for comparison with the Baldwinian-PINNs: 1. the deep architecture consists of similar total number of network weights as the corresponding Baldwinian-PINN models, but distributed across multiple nonlinear hidden layers with smaller number of nodes; 2. the shallow architecture has single nonlinear hidden layer and same number of nodes as the corresponding Baldwinian-PINN model but this necessitates more total network weights because of the additional input variables.
Each architecture also consists of 2 variants: a. tanh activation for the nonlinear hidden layers whereby the network weights are initialized by Xavier method; b. sin activation for the nonlinear hidden layers whereby the network weights are initialized by He method.
This notation is maintained when referencing the respective model performance in Figure 6. For example, the models labelled as DNN-1a and PINN-2b refer to the corresponding baseline DNN model with a deep architecture and tanh activation layer and baseline PINN model with a shallow architecture (single hidden layer) and sin activation layer respectively. Additional model descriptions are in Suppl. LABEL:subsec:baseline.
IV-C1 Direct prediction across tasks with parametric DNNs / PINNs
As a baseline, DNN and PINN models are trained with SGD (ADAM) based on data-driven loss and PINN (data and physics) loss respectively, and applied to predictions for unseen tasks based on interpolation. We explore different MLP architectures and configurations as per Suppl. Table LABEL:tab:baseline_model in order to more fairly compare the best performance across different models.
Critically, when making a direct prediction for a new task, the respective baseline DNN and PINN models do not explicitly incorporate (and guarantee compliance with) the known physics prior. Although it is possible that the parametric PINN has learned a more physically-consistent prediction from the physics loss during training, compliance is not guaranteed for a new task.
In addition, the new tasks must follow an a priori determined input parameterization for interpolation. For example, the DNN and PINN models for convection-diffusion need to be a priori set-up with $(x,\alpha)$ as inputs to enable predictions across different $\alpha$’s. In contrast, the Baldwinian-PINN is parameterization-agnostic, and does not require any input parameterization for predictions of new tasks.
The direct prediction performance of DNNs and PINNs with different model configurations are summarized in Figure 6 and Suppl. Table LABEL:tab:baseline_results. These baseline DNN and PINN predictions have high generalization errors (several orders of magnitude higher than Baldwinian-PINN). As explained above, PINNs do not necessarily outperform DNNs for direct prediction on test tasks (e.g., MSE for convection-diffusion for different model configurations range from $1.7\mathrm{e}{-4}$ to $1.2\mathrm{e}{-2}$ for DNN, and from $1.9\mathrm{e}{-3}$ to $3.4\mathrm{e}{-1}$ for PINN), suggesting that PINN training by itself does not always guarantee generalization in the absence of additional physics-informed fine-tuning.
IV-C2 Prediction across tasks with parametric DNNs / PINNs and physics-informed fine-tuning
In order to assess the utility of a physics-informed fine-tuning step (i.e. the pseudoinverse computation), we optimize the output layer weights of the SGD-trained baseline DNNs and PINNs with a pseudoinverse computation at test time. This fine-tuning procedure is almost identical to the Baldwinian-PINN’s lifetime learning.
The key difference is that the pre-final nonlinear hidden layers weight distribution and learning hyperparameter $\lambda$ in Baldwinian-PINNs are jointly learned by evolution to share among tasks. In contrast, nonlinear hidden layers’ weights for the DNNs and PINNs are SGD-learned from train tasks. As the pseudoinverse solution is highly dependent on $\lambda$, we need to perform a grid search across $\lambda=\{1\mathrm{e}{-2},1\mathrm{e}{-4},1\mathrm{e}{-6},1\mathrm{e}{-8},1\mathrm{e}{-10},1\mathrm{e}{-12},0\}$ to determine the best $\lambda$ during test time and present the solution associated with the lowest LSE. The results after pseudoinverse are compared in Figure 6 and Suppl. Table LABEL:tab:baseline_results.
Results show that physics-informed fine-tuning via pseudoinverse can improve the accuracy of both data-trained DNN and physics-trained PINN models at test time. However, their results are not as generalizable (i.e., the improvements are isolated to certain model and task instances) as the Baldwinian-PINN (which has been optimized for generating accurate solution via pseudoinverse over task distribution).
For example, we observe a significant improvement for the nonlinear kinematics problem (i.e., the best MSE for DNN and PINN improves from $3.0$ to $4.4\mathrm{e}{-5}$ and $1.3$ to $7.9\mathrm{e}{-6}$ respectively after pseudoinverse), potentially because the training distribution is sparser relative to the larger variation in output patterns (see Suppl. LABEL:subsec:additional_ablation).
In contrast, fine-tuning via pseudoinverse doesn’t produce better results than direct prediction for the convection-diffusion problem. The MSE with and without pseudoinverse are $2.3\mathrm{e}{-4}$ and $2.8\mathrm{e}{-2}$, and $1.9\mathrm{e}{-3}$ and $7.2\mathrm{e}{-3}$ for the best-configured DNN and PINN respectively. The pseudoinverse cannot jointly minimize both PDE and IC/BC towards a sufficiently small error, leading to a significantly worse outcome than the direct prediction for some tasks.
While the best DNN model (after pseudoinverse) can be more accurate than the corresponding Baldwinian-PINN on some convection-diffusion tasks at lower $\alpha$, solution quality deteriorates quickly at larger $\alpha$. Interestingly, PINNs with the same configuration (after pseudoinverse) show the opposite trend, highlighting the challenge of obtaining a generalizable model for entire task distributions through SGD, in contrast to Baldwinian neuroevolution.
Overall, these results suggest that an additional step of physics-informed learning (i.e. pseudoinverse) on any new task can be beneficial. However, SGD-trained PINN and DNN models learn pre-final nonlinear hidden layers that are highly variable in their suitability for the physics-informed pseudoinverse-based fine-tuning across new tasks, in contrast to the proposed Baldwinian-PINNs. Hence, these results highlight the advantages of Baldwinian-PINN framework as a complete solution to the challenges encountered in generalizing with physics-informed learning.
V Conclusion
In this paper, we study the Baldwin effect as a novel means to advancing the generalizability of PINNs over a family of governing differential equations. The Baldwin effect is instantiated through a two-stage stochastic programming formulation, wherein the first stage evolves the initial layers of a generalizable PINN model and the second stage trains its final layers to specialize to any new physics task (analogous to lifetime learning). The method is demonstrated to be broadly applicable to the learning of different linear and nonlinear ODE/PDEs encompassing diverse physical phenomena such as convection-diffusion, particle kinematics, and heat and mass transfer. Relative to recent meta-learned PINNs, Baldwinian-PINNs can accelerate the physics-aware predictions by several orders of magnitude, while improving the prediction accuracy by up to one order. A Baldwinian-PINN is thus in the image of a precocial species with accelerated learning ability at birth.
The lifetime learning encapsulated in the Baldwinian paradigm does not require a priori parameterization for the task scenarios, permitting both flexible generalization to new ICs, BCs, and PDE source terms and variations in the lifetime learning task objectives (e.g., different domain and/or sample size). This allows Baldwinian-PINNs to be useful in applications when large number of evaluations with a priori unknown input conditions are sought, e.g. generative design or what-if analysis. In addition, results in Section IV-A2 suggest that Baldwinian-PINNs could be suited to the continual modelling of dynamical systems through their versatility in handling ICs and ability to rapidly model and stitch time windows together with minimal error [51, 52].
In the context of recent interest in foundation models for scientific machine learning [53], specifically through the use of neural operators or neural PDE solvers, our experiments showing accurate and fast generalization across families of linear and nonlinear ODE/PDEs suggest that Baldwinian learning can be an alternate route to such flexible and generalizable machine intelligence models. It will be interesting to test the limits to which Baldwinian-PINNs can learn across broad classes of physics phenomena and/or differential operators in future work.
Lastly, while the experiments in this work focus on optimizing the center and spread of the probability distributions that sample the initial weights in the neural network layers, this can be easily extended to incorporate other state-of-the-art neural architecture search approaches, which directly optimize the graph structure of the node connections. Our experiments indicate that significant improvement relative to other recent meta-learning PINN works can already be observed even when we only optimize the center and spread of the weight sampling distributions under the Baldwinian-PINN framework. The proposed Baldwinian-PINN framework can be seamlessly extended in future work, e.g. via integration with other state-of-the-art neural architecture search algorithms.
Acknowledgment
This research is supported by A*STAR under the AME Programmatic project: Explainable Physics-based AI for Engineering Modelling & Design (ePAI) [Award No. A20H5b0142].